A body of mass 100kg moves at a constant speed of 20 rad / s per circle with a radius of 20cm. How much centripetal force acts on the body?


1
Expert's answer
2021-03-03T11:23:56-0500

"F=ma=m\frac{v^2}{R}=m\frac{(\omega R)^2}{R}=m\omega^2R="


"=100\cdot20^2\cdot0.2=8000\ (N)" .Answer